I need to know how to figure out how fast an object will fall when the weight of the object is known and falling distance of the object is known in inches per second. For example the distance is 8 feet and the weight of the object is 20 pounds. I also need to know what the speed will the object will have reached at the bottom of the 8 foot drop.


1
Expert's answer
2021-12-29T12:34:12-0500

Hello.

The final speed of the falling object barely depends on weight of the object for the given values of height and weight. You can find the final velocity using the formula:


"v_\text{ft/s}=\sqrt{2gh}=\sqrt{2·32·h}=8\sqrt h\text{ (ft/s}).\\
v_\text{in/s}=12v_\text{ft/s}=96\sqrt h\text{ (in/s)}."


However, the height must be converted to feet.

For 8 feet, the speed is 22.6 ft/s or 271.2 in/s.